OpenStudy (anonymous):

39:12 What is the ratio in simplest form of the number of tickets sold I think 4 13

OpenStudy (anonymous):

I meant 4 over 13

OpenStudy (alexandervonhumboldt2):

39/3=13 12/3=4 13:4

OpenStudy (anonymous):

So I was close again

OpenStudy (alexandervonhumboldt2):

can you medal?

OpenStudy (anonymous):

How do I do that? new to this

OpenStudy (alexandervonhumboldt2):

besat responce

OpenStudy (anonymous):

I do that :D

OpenStudy (anonymous):

Can you continue to help me?

OpenStudy (alexandervonhumboldt2):

ys

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

http://prntscr.com/5roxg5 I think it's 2 inch

OpenStudy (alexandervonhumboldt2):

13

OpenStudy (alexandervonhumboldt2):

x=5/25*65